OpenStudy (anonymous):

Is my result for the following integral corect or did I make a mistake? \[\int e ^{-x^{2}}dx\] My result was \[-e^{-2x}/2 + C\]

OpenStudy (anonymous):

I already tried that , erf(x) isn't an accepted result by most teachers :P

OpenStudy (anonymous):

well at least no my teacher :(

OpenStudy (raden):

yeah, you are right but actually, yours still incorrect :P im not sure anyone can solve this

OpenStudy (raden):

no, i cant.. it is so hard :)

OpenStudy (anonymous):

There is no anti-derivative (in terms of elementary functions) for \[e^{-x^2}.\]Whoever gave that to you to solve is playing a cruel joke. If the problem were:\[\int\limits xe^{-x^2}dx\]that would be a different story.

OpenStudy (anonymous):

The best you can hope for (as far as an anti-derivative you can see) is an infinite power series, but most dont except that as an answer.

OpenStudy (anonymous):

thank you

OpenStudy (raden):

welcome, heh :)

OpenStudy (anonymous):

http://www.physicsforums.com/showthread.php?t=127015

OpenStudy (anonymous):

The hard reality of integration is that even after you learn all the different techniques (integration by parts, u-sub, trig-sub, partial fractions, etc), many many many integration problems cannot be solved by hand.

OpenStudy (anonymous):

Maybe it was something like this:\[\int\limits x^3e^xdx\]That can be solved by parts.

OpenStudy (loser66):

ok, you convinced me that it cannot be solved. I give up.

OpenStudy (loser66):

@SithsAndGiggles

OpenStudy (anonymous):

If this were a definite integral, such as over the interval \([0,\infty)\), you could consider \[\left(\int_0^\infty e^{x^2}~dx\right)^2\] then write that as the product of two integrals. Change the variable of one, and you can combine the integrals into a double integral, which you could then work with by converting to polar coordinates. But if this is indeed an indefinite integral... no luck.

OpenStudy (loser66):

ok, got it. I'll be back. thank you

OpenStudy (loser66):

hey friend, you said that you don't know whether the answer is correct or not, I asked my professor about your question, you can trust him since he is a Ph.D in math department in my college and he is willing to answer any of my questions. I am sorry for my ambiguous question lead to his answer without solution, but if you want, I can ask him for that. Here what he said: Dear Professor, This is Hoa. Please, help me with this Can we take integral of e^(-x^2) dx? Thanks in advance You can, because it is continuous and any continuous function can be integrated. You get a function which is more or less the Gaussian probability distribution function. It cannot be expressed in terms of "elementary functions" -- some people will say "you can't integrate it" but they really mean "there is no elementary antiderivative". Reid Huntsinger

OpenStudy (anonymous):

Thanks a lot for your help, after your message I searched the Gaussian probability distribution on google and found this http://en.wikipedia.org/wiki/Gaussian_integral , it solves this exact problem in a few different ways.

OpenStudy (loser66):

that is what SithAndGiggles talked above, isn't it? you bad, me bad!!!ask for help and didn't trust the helper!! have to say sorry to him. "I am sorry, SithAndGiggles""